# Bayesian Controlled FDR Variable Selection via Knockoffs

### v1

@@ 2.3 Controlling the false discovery rate
Moreover, by construction, we have that $\mathbb{P}[W_{j}>0\mid\bm{D},r_{j}=1]>\mathbb{P}[W_{j}<0\mid\bm{D},r_{j}=1]$. Therefore, marginally with respect to $r_{j}$, we obtain
$\mathbb{P}[W_{j}>0\mid\bm{D}]>\mathbb{P}[W_{j}<0\mid\bm{D}].$
By law of total probability, we can write
@@
To complete the proof, we need to show that an invariant prior on $(\beta,\widetilde{\beta})$ leads to a $W$ that satisfies the sign-flip property. This result can be obtained using a similar argument as in Theorem 1 of Gu and Yin, (2021). We report the proof in Appendix A for completeness. ∎
@@
in all applications of this paper. We remark that our model performs variable selection using an upper bound on the posterior probability of non-inclusion. This strategy allows us to overcome issues with strong correlation structure among the covariates, which typically affect the state-of-art variable selection methods that rely on the estimated marginal posterior probabilities of inclusion (see Scenario 3 of our simualtion study below). Our theoretical results, indeed, show that the bound used for our variable selection procedure is not affected by the dependency structure of the covariates.
Although the result in Proposition 1 proves how the method controls the BFDR below a threshold $q$, it assumes that the latent variables $\widetilde{X}$ can be considered as valid knockoffs a priori. As stated by Candès et al., (2018), the construction of valid model-X knockoffs requires perfect knowledge of the covariates’ distribution. However, our proposed model assumes that the covariates $X$ follow a Gaussian distribution where the precision matrix is not known and it has to be estimated. Barber et al., (2020) prove how the knockoff procedure still controls the FDR under a threshold $q$ if the precision matrix of the Gaussian model is correctly estimated. Similarly, we are able to prove that, under mild regularity conditions, the priors chosen to estimate the Gaussian graphical model for the covariates produce a good estimate of the precision matrix. Thus, employing similar arguments as in Barber et al., (2020), we can conclude that, for large sample sizes, our procedure effectively controls the frequentist FDR and, consequently, the expected BFDR over replicated experiments (Guindani et al.,, 2009), even if the precision matrix is unknown. More precisely, we are able to prove that the posterior convergence rate of our estimator is the same as the one of the Bayesian graphical lasso. To show that, we first need to prove that our prior distribution assigns enough probability to values “close” to the true one.
@@
Let the true precision matrix $\Omega_{0}\in\mathcal{U}(\epsilon_{0},l)=\{\Omega:|\{(i,j):1\leq i<j\leq p,\omega_{ij}\neq 0\}|\leq l,0<\epsilon_{0}\leq\text{eig}(\Omega)_{1}<\text{eig}(\Omega)_{p}\leq\epsilon_{0}^{-1}<\infty\}$ for some $0<\epsilon_{0}<\infty$ and $20\leq l\leq p(1-p)/$ 2. Also assume that the prior distributions for the precision matrix $\Omega$ and the adjacency matrix $\bm{G}$ are:
@@
with $v_{g_{ij}}\leq\frac{\|\Omega_{0}\|_{\infty}}{2\log 2}$ and
@@
where $\Pi(\cdot)$ denotes the probability under the prior distribution, $B(p_{\Omega_{0}},\epsilon_{n})=\{p:K(p_{\Omega_{0}},p_{\Omega})\leq\epsilon_{n}^{2},V(p_{\Omega_{0}},p_{\Omega})\leq\epsilon_{n}^{2}\}$ with $K(f,g)=\int f\log(f/g)$ and $V(f,g)=\int f\log^{2}(f/g)$ and $p_{\Omega_{0}}$ and $p_{\Omega}$ are the density function of $N(0,\Omega_{0})$ and $N(0,\Omega)$ respectively.
@@
Following the same argument of Theorem 3.1 in Banerjee and Ghosal, (2015), we can show that
$B(p_{\Omega_{0}},\epsilon_{n})\supset\{p_{\Omega}:\|\Omega-\Omega_{0}\|_{\infty}\leq c\epsilon_{n}/p\}.$
Therefore, it is sufficient to obtain a lower bound for the subset. Moreover, the components of $\Omega$ are not independently distributed, since the prior of $\Omega$ is truncated to the set of positive definite matrices. However, as a small neighborhood of $\Omega_{0}\in\mathcal{U}(\epsilon_{0},l)$ lies within the set of positive definite matrices, the truncation can only increase prior concentration of $B(p_{\Omega_{0}},\epsilon_{n})$. Therefore, we can pretend componentwise independence for the purpose of lower bounding the above prior probability. Then, the probability of interest is
$\Pi(\|\Omega_{0}-\Omega\|_{\infty}\leq c\epsilon_{n}/p)$ $=\Pi(\max_{ij}|\omega_{ij}-\omega_{0,ij}|\leq c\epsilon_{n}/p)$
$=\prod_{ij}\Pi(|\omega_{ij}-\omega_{0,ij}|\leq c\epsilon_{n}/p)$
[…]
×$\times\prod_{i=1}^{p}\int_{c\epsilon_{n}/p-\omega_{0,ii}}^{c\epsilon_{n}/p+\omega_{0,ii}}\frac{\theta}{2}e^{-\frac{\theta}{2}\omega_{ii}}d\omega_{ii}.$
We now focus on the first integral. We obtain
[…]
$\left.+\int_{\omega_{0,ij}-c\epsilon_{n}/p}^{\infty}\frac{1}{\sqrt{2\pi v_{g_{ij}}}}e^{-\frac{\omega_{ij}^{2}}{2v_{g_{ij}}}}d\omega_{ij}\right]$
$\geq 1-\exp\left\{\frac{c\epsilon_{n}/p-\omega_{0,ij}}{2v_{g_{ij}}}\right\}$
$-\exp\left\{\frac{-c\epsilon_{n}/p-\omega_{0,ij}}{2v_{g_{ij}}}\right\},$
where the bound is obtained by Chernoff inequality. We need to show that
$\lim_{\epsilon_{n}\to 0}\frac{c\epsilon_{n}/p}{1-\exp\left\{\frac{c\epsilon_{n}/p-\omega_{0,ij}}{2v_{g_{ij}}}\right\}-\exp\left\{\frac{-c\epsilon_{n}/p-\omega_{0,ij}}{2v_{g_{ij}}}\right\}}=0^{+},$
since this would imply
$\int_{c\epsilon_{n}/p-\omega_{0,ij}}^{c\epsilon_{n}/p+\omega_{0,ij}}$ $\frac{1}{\sqrt{2\pi v_{g_{ij}}}}e^{-\frac{\omega_{ij}^{2}}{2v_{g_{ij}}}}d\omega_{ij}$
$\geq 1-\exp\left\{\frac{c\epsilon_{n}/p-\omega_{0,ij}}{2v_{g_{ij}}}\right\}-\exp\left\{\frac{-c\epsilon_{n}/p-\omega_{0,ij}}{2v_{g_{ij}}}\right\}$
$\gtrsim c\epsilon_{n}/p.$
Noting that $1-\exp\left\{\frac{c\epsilon_{n}/p-\omega_{0,ij}}{2v_{g_{ij}}}\right\}-\exp\left\{\frac{-c\epsilon_{n}/p-\omega_{0,ij}}{2v_{g_{ij}}}\right\}$ is positive when $\epsilon_{n}=0$ under the conditions for $v_{g_{ij}}$ is sufficient to obtain the limit result.
Focusing now on the term referring to diagonal elements of $\Omega$, we have that
[…]
$-\exp\left\{-\frac{\theta}{2}(c\epsilon_{n}/p+\omega_{0,ii})\right\}.$
Using the same argument as above, we can show
[…]
since the limit of the same ratio as previous term for $\epsilon_{n}$ to 00 converges to zero. These two results complete the proof. ∎
Using the prior concentration result stated in Proposition 2, we can investigate the posterior convergence rate of the graphical parameters. Proposition 3 shows how the posterior convergence rate of our model is equal to the convergence rate of the graphical lasso (Friedman et al.,, 2008) and the Bayesian graphical lasso (Park and Casella,, 2008). Note that, in Proposition 3, the prior distribution for the precision matrix in Equation (19) is slightly different from the one used in our model in Equation (11) since it includes the probability $P(\bar{R}\geq|\bm{G}|)$ to penalize denser graphs. However, in the context of MCMC algorithm to sample from posterior distribution, such probability has little to no effect. The ratio of this probabilities is, indeed, close or superior to 1 for most proposal distributions.
@@
Let $X^{(n)}=(X_{1},\dots,X_{n})$ be a random sample from the p-dimensional Gaussian distribution with mean 00 and precision matrix $\Omega_{0}\in\mathcal{U}(\epsilon_{0},l)=\{\Omega:|\{(i,j):1\leq i<j\leq p,\omega_{ij}\neq 0\}|\leq l,0<\epsilon_{0}\leq\text{eig}(\Omega)_{1}<\text{eig}(\Omega)_{p}\leq\epsilon_{0}^{-1}<\infty\}$ for some $0<\epsilon_{0}<\infty$ and $20\leq l\leq p(1-p)/$ 2. Also assume that the prior distributions are:
@@
with $v_{g_{ij}}\leq\frac{\|\Omega_{0}\|_{\infty}}{2\log 2}$ and
$p(\bm{G}\mid v_{0},v_{1},\theta,\xi)=\frac{K(\bm{G},v_{0},v_{1},\theta)}{K(v_{0},v_{1},\xi)}\prod_{j<j^{\prime}}\xi^{g_{jj^{\prime}}}(1-\xi)^{1-g_{jj^{\prime}}}P(\bar{R}\geq|\bm{G}|),$ (19)
with $\xi<0.5$ and $P(\bar{R}>a_{1}m)\leq\exp\{-a_{2}m\log m\}$ for some constants $a_{1}$ and $a_{2}$ and $m=1,2,\dots$. Then,
$\mathbb{E}_{0}[P\{\|\Omega-\Omega_{0}\|_{2}>M\epsilon_{n}\mid X^{(n)}\}]\to 0$
@@
Similarly to Banerjee and Ghosal, (2015), this proof uses the general theory of posterior convergence rate developed in Ghosal et al., (2000). First, we need to prove that the prior probability assigns enough mass to the Kullback-Leibler neighborhood of the true value. Then we show that the space in which the true parameter is contained is not too big in terms of metric entropy with respect to the Frobenius distance. Finally, we prove that the prior probability for matrices not contained in the the same space of the true value is not too high.
More in detail, Proposition 2 shows that
$\Pi(B(p_{\Omega_{0}},\epsilon_{n}))\gtrsim(c\epsilon_{n}/p)^{p+l}$
where $B(p_{\Omega_{0}},\epsilon_{n})=\{p:K(p_{\Omega_{0}},p_{\Omega})\leq\epsilon_{n}^{2},V(p_{\Omega_{0}},p_{\Omega})\leq\epsilon_{n}^{2}\}$ with $K(f,g)=\int f\log(f/g)$ and $V(f,g)=\int f\log^{2}(f/g)$. By Theorem 2.1 of Ghosal et al., (2000), this implies
$(p+l)(\log p+\log\epsilon_{n}^{-1})\asymp n\epsilon_{n}^{2}$
so as to get $\epsilon_{n}=n^{-1/2}(p+l)^{1/2}(\log p)^{1/2}$ as the prior concentration rate around $\Omega_{0}$. Using the same argument of Theorem 3.1 in Banerjee and Ghosal, (2015), we can show that the metric entropy of a subset of $\mathcal{U}$ is bounded. In particular, consider the set $\mathcal{P}_{n}$ to be the space of all densities $p_{\Omega}$ such that the graph induced by $\Omega$ has maximum number of edges $\bar{r}<p(p-1)/4$ and each entry of $\Omega$ is at most $L$ in absolute value, where $\bar{r}$ and $L$ depend on $n$ and are to be chosen later. Then, the metric entropy of the set of precision matrices with respect to the Frobenius distance is given by
$\log\left\{\sum_{j=1}^{\bar{r}}\left(\frac{L}{\epsilon_{n}}\right)^{j}\binom{\binom{p}{2}}{j}\right\}$ $\leq\log\left\{\bar{r}\left(\frac{L}{\epsilon_{n}}\right)^{\bar{r}}\binom{p+\binom{p}{2}}{\bar{r}}\right\}$
$\lesssim\log\bar{r}+\bar{r}\log L+\bar{r}\log\epsilon_{n}^{-1}+\bar{r}\log p,$
so the choices $\bar{r}\sim b_{1}n\epsilon_{n}^{2}/\log n$, and $L\sim b_{2}n\epsilon_{n}^{2}$ for any choice of the constants $b_{1}$ and $b_{2}>0$ will satisfy the rate equation
$\log\bar{r}+\bar{r}\log p+\bar{r}\log\epsilon_{n}^{-1}+\bar{r}\log(n\epsilon_{n}^{2})\asymp n\epsilon_{n}^{2}.$
Note that, under condition $n\epsilon_{n}^{2}/\log n\ll p(p-1)/2$, the requirement $\bar{r}<p(p-1)/4$ is satisfied as $n\to\infty$.
Finally, we need to bound the prior probability of the complementary set $\mathcal{P}_{n}^{c}$. Note that a distribution $p_{\Omega}$ can be outside of $\mathcal{P}_{n}$ only if the graph induced by $\Omega$ has more edges than $\bar{r}$ or if any entry of the matrix $\Omega$ has larger magnitude than $L$. The probability of the second event is bounded by $p(1-p)/2\,e^{-L^{2}}$ which, by the choice of $L\sim b_{2}n\epsilon_{n}^{2}$, is bounded by $\exp\{-b_{3}n\epsilon_{n}\}$ for some constant $b_{3}$. Since $b_{2}$ can be taken as large as we want, $b_{3}$ can also be taken as large as required. The prior probability of having more than $\bar{r}$ edges is
$P(\bar{R}>\bar{r})\leq\exp(-a^{\prime}_{2}b_{1}n\epsilon_{n}^{2}),$
where $a^{\prime}_{2}b_{1}$ can be made as large as possible by making $b_{1}$ large. Hence $\epsilon_{n}=n^{-1/2}(p+l)^{1/2}(\log p)^{1/2}$ is the posterior convergence rate. ∎
Proposition 3 shows that, with sufficiently small prior variances, the posterior mean of the precision matrix $\Omega$ converges to the true one with the same rate as the convergence rate of graphical lasso (Rothman et al.,, 2008). Therefore, employing similar arguments as in Barber et al., (2020), for a sufficiently large sample size, our model controls the frequentist FDR and, consequently, the expected BFDR under a threshold $q$ without the knowledge of the true precision matrix $\Omega$.
@@ 3.3 Results and comparisons
Let us first examine a single data set to gain insight into how the proposed model works. Specifically, we focus on a randomly selected data set generated under Scenario 2. Figure 2 shows the posterior distribution of the feature statistics $W_{j}$, $j=1,\ldots,p$, and Figure 3 shows the behavior of the proposed Bayesian knockoffs on this simulated data set. Bars in this plot correspond to quantities ^$\widehat{\mathbb{P}}[r_{j}=0\mid\bm{D}]$, $j=1,\ldots,p$, calculated as in Equation (18), while points represent the estimated $\text{BFDR}(\mathcal{S})$ with $\mathcal{S}$ the set of all previous indexes. The dashed red line indicates the chosen threshold $q=0.1$. Points in red correspond to the true active variables and those in green to the inactive ones. The final set of selected variables is ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{\mathcal{S}}=}\{X_{1},X_{2},X_{6},X_{24},X_{30}\}$ and it contains exactly the true active variables, showing that the model can effectively distinguish between active and inactive variables.
Figure 2: Simulation study: Posterior distribution of $W_{j}$, $j=1,\ldots,p$, for a simulated data set under Scenario 2. Those highlighted in blue are selected according to Figure 3.
Figure 3: Simulation study: Behavior of the proposed method for variable selection for a simulated data set under Scenario 2. Bars represent estimates ^$\widehat{\mathbb{P}}[r_{j}=0\mid\bm{D}]$, $j=1,\ldots p$, in increasing order and points represent the estimated $\text{BFDR}(\mathcal{S})$ with $\mathcal{S}$ the set of all previous indexes. The dashed red line is the chosen threshold $q=0.1$. Points in red correspond to true active variables and those in green to inactive variables.
@@
Next, we report the comparison of the performances of our proposed model with other procedures for variable selection with and without FDR control. Among these procedures, we include the classical model-X knockoff filter proposed by Barber and Candès, (2019). In particular, we consider two versions of model-X knockoff filter, one with the true covariance matrix assumed known, and one with an estimate of $\bm{\Sigma}$ obtained via the graphical lasso estimator of Friedman et al., (2008). Indeed, Barber et al., (2020) proved that, under certain conditions, the graphical lasso can be used to preserve the FDR guarantees of knockoff filters. For both our model and the classical knockoff filter, we fix the FDR threshold at $q=0.1$. The knockoffs were generated using the R package knockoffs, and the covariance matrix was estimated using the glasso package. In addition to the model-X knockoff filter, our comparison includes other state-of-art methods for variable selection. In particular, we include the Lasso regression (Tibshirani,, 1996) and a Bayesian linear regression model with discrete spike-and-slab prior and a Beta hyperprior on the inclusion probability. We also considered the joint edge and regression estimation model proposed by Peterson et al., (2016), which we implemented ourselves as a version of our proposed model without knockoffs latent variables. The regularizing parameter for lasso regression was determined by minimizing the cross-validated error using the R package glmnet. We note that this method does not employ any FDR correction. Variable selection for the Bayesian and joint estimation methods was achieved by thresholding the marginal posterior probabilities of inclusion (PPIs) at $0.5$, following the median model approach (Barbieri and Berger,, 2004). We also considered the BFDR approach proposed by Newton et al., (2004).
@@
Results in Table 1 indicate that the proposed Bayesian method succeeds in keeping the FDR below the given threshold in every scenario, while achieving strong performances in TPR compared to the other methods, as well as MCC and F1 scores. It is worth noting that the classical knockoff filter exhibits low TPR in Scenarios $1$ and $2$, where the magnitude of the true coefficients is rather small. In other settings, not reported here, where the signal was strong and the sample size larger, we noticed improved performances for all methods, including classical knockoffs. Further, the values of the standard deviations for the TPR of the knockoff filter in the first two scenarios are higher than those of the other methods. This happens because, across the $100$ simulated data sets, the knockoff filter TPR is often zero and sometimes quite high. Finally, we note that classical knockoff has a slightly higher FDR than the threshold $0.1$ when the covariates are not independent and the covariance matrix is estimated using graphical lasso. As shown in Barber et al., (2020), the classical knockoff filter does not control FDR exactly below $q$ when the covariance matrix used to build the knockoffs is not the true one but an estimate. In our case, the graphical lasso tends to produce a sparser than needed graph. This behavior also supports our choice of setting the parameter of the priors in the proposed method to obtain a generally dense graph to be used to generate the knockoffs. Results for the third scenario confirm that the proposed Bayesian Knockoff filter is also able to outperform the state-of-art methods in settings with larger dimensionality. As a matter of fact, it successfully controls FDR under $0.1$, unlike the lasso regression. Moreover, the TPR is on average comparable to the one of the classical knockoff filter but it is more stable as it has a much lower standard deviation.
@@
3.4 Accelerated Failure Time models

### v2

@@ 2.3 Controlling the false discovery rate
By the law of total probability, we can write
@@
To complete the proof, we need to show that an invariant prior on $(\beta,\widetilde{\beta})$ leads to a $W$ that satisfies the sign-flip property. This result can be obtained using a similar argument as in Theorem 1 of Gu and Yin, (2021). We report the proof in the Supplementary Material for completeness. ∎
@@
in all applications of this paper. We remark that our model performs variable selection using an upper bound on the posterior probability of non-inclusion. This strategy allows us to overcome issues with strong correlation structure among the covariates, which typically affect the state-of-art variable selection methods that rely on the estimated marginal posterior probabilities of inclusion, as empirically shown in our simulation study. Our theoretical results, indeed, show that the bound used for our variable selection procedure is not affected by the dependency structure of the covariates.
Although Proposition 1 shows that our method controls the BFDR at a threshold $q$, it assumes that the latent variables $\widetilde{X}$ are valid knockoffs a priori. As discussed by Candès et al., (2018), the construction of valid model-X knockoffs requires perfect knowledge of the covariates’ distribution. However, our proposed model assumes that the covariates $X$ follow a Gaussian distribution, with an unknown precision matrix that must be estimated from the data. Barber et al., (2020) prove that the knockoff procedure maintains FDR control at a desired threshold $q$ if the precision matrix of the Gaussian model is correctly estimated. Here, we show that, under mild regularity conditions, the priors chosen to estimate the Gaussian graphical model for the covariates yield accurate posterior estimates of the precision matrix. Thus, drawing on similar arguments as in Barber et al., (2020), we can conclude that our procedure controls both the frequentist FDR and, consequently, the expected BFDR (Guindani et al.,, 2009) in large samples, even if the precision matrix is unknown. More precisely, we are able to prove that the posterior convergence rate of our estimator is the same as the one of the Bayesian graphical lasso. To show such result, we first need to prove that our prior distribution assigns enough probability mass to values “close” to the true one.
@@
Let the true precision matrix ${\bm{\Omega}}_{0}\in\mathcal{U}(\epsilon_{0},l)=\{{\bm{\Omega}}:|\{(i,j):1\leq i<j\leq p,\omega_{ij}\neq 0\}|\leq l,0<\epsilon_{0}\leq\text{eig}({\bm{\Omega}})_{1}<\text{eig}({\bm{\Omega}})_{p}\leq\epsilon_{0}^{-1}<\infty\}$ for some $0<\epsilon_{0}<\infty$ and $20\leq l\leq p(1-p)/$ 2. Also assume that the prior distributions for the precision matrix $\Omega$ and the adjacency matrix $\bm{G}$ are:
@@
with $v_{g_{ij}}\leq\frac{\|{\bm{\Omega}}_{0}\|_{\infty}}{2\log 2}$ and
@@
with $\Pi(\cdot)$ the probability under the prior distribution, $B(p_{\Omega_{0}},\epsilon_{n})=\{p:K(p_{\Omega_{0}},p_{\Omega})\leq\epsilon_{n}^{2},V(p_{\Omega_{0}},p_{\Omega})\leq\epsilon_{n}^{2}\}$ with $K(f,g)=\int f\log(f/g)$ and $V(f,g)=\int f\log^{2}(f/g)$ and $p_{\Omega_{0}}$ and $p_{\Omega}$ the density functions of $N(0,{\bm{\Omega}}_{0})$ and $N(0,{\bm{\Omega}})$ respectively.
@@
See the Supplementary Material. ∎
Using the prior concentration result stated in Proposition 2, we can investigate the posterior convergence rate of the graphical parameters. Proposition 3 shows how the posterior convergence rate of our model is equal to the convergence rate of the graphical lasso (Friedman et al.,, 2008) and the Bayesian graphical lasso (Park and Casella,, 2008). Note that, in Proposition 3, the prior distribution for the precision matrix in Equation (19) is slightly different from the one used in our model in Equation (11) since it includes the probability $\mathbb{P}[\bar{R}\geq|\bm{G}|]$ to penalize denser graphs. However, in the context of MCMC algorithm to sample from posterior distribution, such probability has little to no effect. The ratio of these probabilities is, indeed, close or superior to 1 for most proposal distributions.
@@
Let $X^{(n)}=(X_{1},\dots,X_{n})$ be a random sample from the p-dimensional Gaussian distribution with mean 00 and precision matrix ${\bm{\Omega}}_{0}\in\mathcal{U}(\epsilon_{0},l)=\{{\bm{\Omega}}:|\{(i,j):1\leq i<j\leq p,\omega_{ij}\neq 0\}|\leq l,0<\epsilon_{0}\leq\text{eig}({\bm{\Omega}})_{1}<\text{eig}({\bm{\Omega}})_{p}\leq\epsilon_{0}^{-1}<\infty\}$ for some $0<\epsilon_{0}<\infty$ and $20\leq l\leq p(1-p)/$ 2. Also assume that the prior distributions are:
@@
with $v_{g_{ij}}\leq\frac{\|{\bm{\Omega}}_{0}\|_{\infty}}{2\log 2}$ and
$p(\bm{G}\mid v_{0},v_{1},\theta,\xi)=\frac{K(\bm{G},v_{0},v_{1},\theta)}{K(v_{0},v_{1},\xi)}\prod_{j<j^{\prime}}\xi^{g_{jj^{\prime}}}(1-\xi)^{1-g_{jj^{\prime}}}\mathbb{P}(\bar{R}\geq|\bm{G}|),$ (19)
with $\xi<0.5$ and $\mathbb{P}(\bar{R}>a_{1}m)\leq\exp\{-a_{2}m\log m\}$ for some constants $a_{1}$ and $a_{2}$ and $m=1,2,\dots$. Then,
$\mathbb{E}_{0}[\mathbb{P}\{\|{\bm{\Omega}}-{\bm{\Omega}}_{0}\|_{2}>M\epsilon_{n}\mid X^{(n)}\}]\to 0$
@@
See the Supplementary Material. ∎
Proposition 3 shows that, with sufficiently small prior variances, the posterior mean of the precision matrix ${\bm{\Omega}}$ converges to the true one with the same rate as the convergence rate of graphical lasso (Rothman et al.,, 2008). Therefore, employing similar arguments as in Barber et al., (2020), for a sufficiently large sample size, our model controls the frequentist FDR and, consequently, the expected BFDR under a threshold $q$ without the knowledge of the true precision matrix ${\bm{\Omega}}$.
@@ 3.3 Results and comparisons
Let us first examine a single data set to gain insight into how the proposed model works. Specifically, we focus on a randomly selected data set generated under Scenario 3. The top plot of Figure 2 shows the estimated BFDR and the upper bound for $\mathbb{P}[r_{j}=0\mid\bm{D}]$, for $j=1,\ldots p$ and covariates sorted by the BFDR values. Our method correctly identifies 18 of the 24 true predictors with 6 false negatives, for a $q=0.1$ threshold on the BFDR. The bottom plot of Figure 2 reports the PPIs of all covariates, calculated by fitting a traditional Bayesian linear regression model with discrete spike-and-slab priors and a Beta hyperprior on the inclusion probability of the latent indicators (George and McCulloch,, 1997; Brown et al.,, 1998; Vannucci,, 2021). This method selects 21 of the 24 true predictors with 44 false positives, for a $0.5$ threshold on the PPIs. The large number of false positives is due to the fact that the presence of collinearity in the data greatly affects the PPIs values, resulting in inflated values for inactive covariates that are correlated to the active ones. This simple example highlights the inability of clearly dividing active and inactive variables based on the posterior probability values. The results are confirmed in the comparison study below, that shows performance values averaged over replicated datasets. This empirical comparison of our Bayesian Knockoff with a traditional spike and slab regression confirms the theoretical results in Theorem 1, where the upper bound is unaffected by the covariates’ dependency structure.
Figure 2: Simulation study: Variable selection performed by our Bayesian knockoff filter method (BayeKnock) and by spike-and-regression (Spike-Slab) on a simulated dataset under Scenario 3. Top plot (BayesKnock): Bars represent estimates of the upper bound for ${\mathbb{P}}[r_{j}=0\mid D]$ as in Equation (17) while points represent the estimated BFDR. Red points correspond to true active variables. Bottom plot (Spike-Slab): Estimated marginal posterior probabilities of inclusion (PPIs). Red bars correspond to true active variables.
@@
Next, we report results of our method averaged over 100 replicated datasets and perform a comparison of the performances with other procedures for variable selection, with and without FDR control. Among these procedures, we include the classical model-X knockoff filter proposed by Barber and Candès, (2019). In particular, we consider two versions of the model-X knockoff filter, one with the true covariance matrix assumed known, and one with an estimate of $\bm{\Sigma}$ obtained via the graphical lasso estimator of Friedman et al., (2008). Indeed, Barber et al., (2020) proved that, under certain conditions, the graphical lasso can be used to preserve the FDR guarantees of knockoff filters. For both our model and the classical knockoff filter, we fix the FDR threshold at $q=0.1$. The knockoffs were generated using the R package knockoffs, and the covariance matrix was estimated using the glasso package. In addition to the model-X knockoff filter, our comparison includes other state-of-art methods for variable selection. In particular, we include the Lasso regression (Tibshirani,, 1996) and a Bayesian linear regression model with discrete spike-and-slab prior and a Beta hyperprior on the inclusion probability, which we implemented ourselves. We also considered the joint edge and regression estimation model proposed by Peterson et al., (2016), which we implemented ourselves as a version of our proposed model without knockoffs latent variables. The regularizing parameter for lasso regression was determined by minimizing the cross-validated error using the R package glmnet. We note that this method does not employ any FDR correction. Variable selection for the Bayesian and joint estimation methods was achieved by thresholding the marginal posterior probabilities of inclusion (PPIs) at $0.5$, following the median model approach (Barbieri and Berger,, 2004). We also considered the BFDR approach proposed by Newton et al., (2004).
@@
Results in Table 1 indicate that the proposed Bayesian method succeeds in keeping the FDR below the given threshold in every scenario, while achieving strong performances in TPR compared to the other methods, as well as MCC and F1 scores. It is worth noting that the classical knockoff filter exhibits low TPR in Scenarios $1$ and $2$, where the magnitude of the true coefficients is rather small. In other settings, not reported here, where the signal was strong and the sample size larger, we noticed improved performances for all methods, including classical knockoffs. Further, the values of the standard deviations for the TPR of the knockoff filter in the first two scenarios are higher than those of the other methods. This happens because, across the $100$ simulated data sets, the knockoff filter TPR is often zero and sometimes quite high. Finally, we note that classical knockoff has a slightly higher FDR than the threshold $0.1$ when the covariates are not independent and the covariance matrix is estimated using graphical lasso. As shown in Barber et al., (2020), the classical knockoff filter does not control FDR exactly below $q$ when the covariance matrix used to build the knockoffs is not the true one but an estimate. In our case, the graphical lasso tends to produce a sparser than needed graph. This behavior also supports our choice of setting the parameter of the priors in the proposed method to obtain a generally dense graph to be used to generate the knockoffs. Results for the third scenario confirm that the proposed Bayesian Knockoff filter is also able to outperform the state-of-art methods in settings with larger dimensionality. As a matter of fact, it successfully controls FDR under $0.1$, unlike the lasso regression. Moreover, the TPR is on average comparable to the one of the classical knockoff filter, but it is more stable as it has a much lower standard deviation.
@@
3.4 Simulation of a survival outcome from an accelerated failure time model